I was doing great until I ran into this one please help!
Find all positive values for k for which each of the following can be factored
x^2 + 5x + k
Thank you very much!

Answer by edjones(8007)

x^2 + 5x + k
k=4 (4,1)
k=6 (3,2)
k=25/4 (5/2,5/2) perfect square.
There are an infinate number of fractions, though. eg. k=5.25 (1.5,3.5)
Ed